# Marginal treatment effects in the absence of instrumental variables

### v2

@@ Appendix A Proof of Theorem 1
$\partial_{k}m_{d0}\left(x^{C}\right)\partial_{j}\pi_{0}\left(x^{C}\right)-\partial_{j}m_{d0}\left(x^{C}\right)\partial_{k}\pi_{0}\left(x^{C}\right)=\partial_{j}\pi_{0}\left(x^{C}\right)\beta_{d,k}^{C}-\partial_{k}\pi_{0}\left(x^{C}\right)\beta_{d,j}^{C},$ (A.2)
@@
$\left(\begin{array}[]{c}\partial_{k}m_{d0}\left(x^{C}\right)\partial_{j}\pi_{0}\left(x^{C}\right)-\partial_{j}m_{d0}\left(x^{C}\right)\partial_{k}\pi_{0}\left(x^{C}\right)\\ \partial_{k}m_{d0}\left(\tilde{x}^{C}\right)\partial_{j}\pi_{0}\left(\tilde{x}^{C}\right)-\partial_{j}m_{d0}\left(\tilde{x}^{C}\right)\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\end{array}\right)=\Psi\left(\begin{array}[]{c}\beta_{d,k}^{C}\\ \beta_{d,j}^{C}\end{array}\right),$ (A.3)
@@
$\Psi=\left(\begin{array}[]{cc}\partial_{j}\pi_{0}\left(x^{C}\right)&-\partial_{k}\pi_{0}\left(x^{C}\right)\\ \partial_{j}\pi_{0}\left(\tilde{x}^{C}\right)&-\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\end{array}\right).$
Assumption NL2.(v) ensures that the determinant of $\Psi$ is nonzero, which shows that $\Psi$ is nonsingular. Therefore, equation (A.3) can be solved for $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$ by inverting $\Psi$, thereby identifying $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. Given identification of $\beta_{d,k}^{C}$, we can then identify all other coefficient $\beta_{d,l}^{C}$ in $\beta_{d}^{C}$ by solving (A.2) with the subscript $j$ replaced by $l$, which gives
@@
Appendix B Further discussion on the nonlinearity assumption
@@
When $\dim\left(X^{C}\right)=1$, it follows from (A.1) and the assumption of Theorem B.1 that
@@
The remaining part of the proof is the same as the proof of Theorem 1.
The assumption of Theorem B.1 in place of NL1 requires the univariate function $\pi_{0}$ to be not one-to-one, but imposes no other smoothness or continuity restriction on $\pi_{0}$. Therefore, it doesn’t require $X^{C}$ to be continuously valued. In the extreme case of $X^{C}$ containing only one binary covariate, this assumption will be equivalent to the full irrelevance of $X^{C}$ to the treatment probability, which is a condition suggested by Chamberlain (1986) for identification of the sample selection model. The following two theorems show that Assumption NL1 and the assumption of Theorem B.1 can be extended to the case of multiple continuous covariates.
Theorem B.2.
@@
It follows from equation (A.1) and $\pi_{0}\left(x^{C}\right)=\pi_{0}\left(\tilde{x}^{C}\right)$ that
@@
$\partial_{k}m_{d0}\left(\breve{x}^{C}\right)$ $=$ $\beta_{d,k}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)\partial_{k}\pi_{0}\left(\breve{x}^{C}\right),$ (B.1)
$\partial_{j}m_{d0}\left(\breve{x}^{C}\right)$ $=$ $\beta_{d,j}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)\partial_{j}\pi_{0}\left(\breve{x}^{C}\right).$ (B.2)
Since $\partial_{k}\pi_{0}\left(\breve{x}^{C}\right)\neq 0$ by condition (ii), we can find $g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)$ from (B.1) to be
@@
and substitute it into (B.2) to identify $\beta_{d,j}^{C}$ as
@@
Therefore, $\beta_{d}^{C}$ is identified. The remaining part of the proof is the same as the proof of Theorem 1.
Theorem B.3.
@@
The identification of $\beta_{d,k}^{C}$ is analogous to Theorem 1. Then, the identification of all the other coefficients $\beta_{d,j}^{C}$ in $\beta_{d}^{C}$ and the remaining components of the model is the same as in Theorem B.2 and Theorem 1, respectively.
@@
For each $k\in\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$, taking the partial derivative of both sides of (A.1) with respect to $x_{k}^{C}$ gives
@@
which immediately identifies all $\beta_{d,k}^{C}$ in $\beta_{d}^{C}$. The remaining part of the proof is the same as that of Theorem 1.
The above discussion further illustrates the overidentification feature of our strategy. In fact, the identifying power of our method is determined by the number of nonlinearity points in the data satisfying Assumption NL or Theorems B.1-B.4, given different benchmark values of $X^{D}$.
Appendix C Limited valued outcomes
@@
Assumption C.L (Linear index). Assume that $Y_{d}=l\left(X^{\prime}\gamma_{d},V_{d}\right)$, $d=0,1$, holds for some known link function $l$ and unknown coefficients $\gamma_{d}$, where $V_{d}$ is unobserved error term of $Y_{d}$.
@@
Assumption C.N (Normalization). Decompose $X^{\prime}\gamma_{d}=X^{C\prime}\gamma_{d}^{C}+X^{D\prime}\gamma_{d}^{D}$, where $X^{C}$ and $X^{D}$ consist of covariates that are continuously and discretely distributed, respectively. Assume that $\gamma_{d,1}^{C}$, the first element of $\gamma_{d}^{C}$, equal to 1.
Assumption C.N imposes the convenient normalization that the first continuous covariate has a unit coefficient. This scaling of $\gamma_{d}$ is arbitrary and is innocuous because our focus is on the identification of MTE, which is a difference in the conditional expectations of $l\left(X^{\prime}\gamma_{d},V_{d}\right)$, rather than on separate identification of $\gamma_{d}$ and $l$. Since in the generalized linear model the potential outcomes are not necessarily additive in the error term, the mean independence assumption needs to be strengthened to a stricter distributional independence assumption.
Assumption C.CI (Conditional Independence). Assume that $\left.V_{d}\perp\!\!\!\!\perp X\right|V$ for $d=0,1$, namely, $V_{d}$ is independent of $X$ conditional on $V$, where $V$ is the reduced-form treatment error in equation (3).
Recalling that $V\perp\!\!\!\!\perp X$ by definition, Assumption C.CI is equivalent to the full independence $\left(V_{d},V\right)\perp\!\!\!\!\perp X$ for $d=0,1$, which implies that both the marginal distribution of $V_{d}$ and the copula of $V_{d}$ and $V$ are independent of $X$. Under Assumptions C.L and C.CI, we have
@@
The additive nonseparability of the observables and unobservables constitutes the primary difficulty in identifying MTE in the limited outcome case. Meanwhile, Assumptions C.L and C.CI lead to a double index form of the observable outcome regression functions for each treatment status, that is,
$E\left[\left.Y1\left\{D=d\right\}\right|X=x\right]=r_{d}\left(x^{\prime}\gamma_{d},\pi\left(x\right)\right)$ (C.1)
@@
[…]
Provided that $\pi\left(x\right)$ is a nonlinear index, $\gamma_{d}$ and thus $r_{d}$ can be identified based on functional form, which ensures identification of MTE since
$\Delta^{\text{MTE}}\left(x,v\right)=\partial_{2}r_{1}\left(x^{\prime}\gamma_{1},v\right)+\partial_{2}r_{0}\left(x^{\prime}\gamma_{0},v\right),$ (C.4)
@@
Assumption C.NL (Non-Linearity). Assume that the functions $\pi_{0}$ and $r_{d}$ are differentiable and denote their partial derivatives with respect to the $k$-th argument as $\partial_{k}\pi_{0}$ and $\partial_{k}r_{d}$, where $\pi_{0}\left(x^{C}\right)=\pi\left(x^{C},0\right)$ and $r_{d}$, $d=0,1$, are defined in (C.1) and (C.1). Assume that there exist two vectors $x^{C}$ and $\tilde{x}^{C}$ on the support of $X^{C}$ and two elements $k$ and $j$ of the set $\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$ such that (i) $\partial_{1}r_{d}\left(x^{C\prime}\gamma_{d}^{C},\pi_{0}\left(x^{C}\right)\right)\neq 0$, (ii) $\partial_{1}r_{d}\left(\tilde{x}^{C\prime}\gamma_{d}^{C},\pi_{0}\left(\tilde{x}^{C}\right)\right)\neq 0$, (iii) $\partial_{k}\pi_{0}\left(x^{C}\right)\neq\partial_{1}\pi_{0}\left(x^{C}\right)\gamma_{d,k}^{C}$, (iv) $\partial_{j}\pi_{0}\left(x^{C}\right)\neq\partial_{1}\pi_{0}\left(x^{C}\right)\gamma_{d,j}^{C}$, (v) $\partial_{l}\pi_{0}\left(\tilde{x}^{C}\right)\neq\partial_{1}\pi_{0}\left(\tilde{x}^{C}\right)\gamma_{d,l}^{C}$ for $l=2,\cdots,\dim\left(X^{C}\right)$, and
@@
Assumption C.NL.(i)-(ii) require $r_{d}$ to depend on the linear index, and (iii)-(vi) essentially require some nonlinear variation in $\pi_{0}$ under the scale normalization. As in Assumption NL, it is difficult to construct examples other than $\pi_{0}\left(x^{C}\right)=f\left(x^{C\prime}\gamma\right)$ that violates Assumption C.NL. Finally, a support assumption and an invertibility assumption are imposed for technical reasons.
Assumption C.S (Support). For each $k\in\left\{1,2,\cdots,\dim\left(X^{D}\right)\right\}$, assume for some $x_{k}^{D}\neq 0$ in the support of $X_{k}^{D}$ that there exists $x^{C}\left(k\right)$ in the support of $X^{C}$ such that $\pi\left(x^{C}\left(k\right),x^{Dk}\right)$ is in the support of $\pi_{0}\left(X^{C}\right)$ and that $x^{C}\left(k\right)^{\prime}\gamma_{d}^{C}+x_{k}^{D}\gamma_{d,k}^{D}$ is in the support of $X^{C\prime}\gamma_{d}^{C}$ for $d=0,1$.
Assumption C.I (Invertibility). Assume that the function $r_{d}$ is invertible on its first argument for $d=0,1$.
[…]
Theorem C.1.
If Assumptions C.L, C.NL, C.CI, C.S, C.N, and C.I hold, then $\gamma_{d}$ and $r_{d}\left(t,p\right)$ at all points in the support of $X^{\prime}\gamma_{d}$ and $P$ are identified for $d=0,1$.
Having established identification of $\gamma_{d}$ and $r_{d}$, the MTE for limited valued outcomes can be identified by (C.4) and estimated by the semiparametric least squares method of Escanciano et al. (2016). Alternatively, the LIV estimation procedure may be adapted according to

### v3

@@ Appendix B Proof of Theorem 1
$\partial_{k}m_{d0}\left(x^{C}\right)\partial_{j}\pi_{0}\left(x^{C}\right)-\partial_{j}m_{d0}\left(x^{C}\right)\partial_{k}\pi_{0}\left(x^{C}\right)=\partial_{j}\pi_{0}\left(x^{C}\right)\beta_{d,k}^{C}-\partial_{k}\pi_{0}\left(x^{C}\right)\beta_{d,j}^{C},$ (B.2)
@@
$\left(\begin{array}[]{c}\partial_{k}m_{d0}\left(x^{C}\right)\partial_{j}\pi_{0}\left(x^{C}\right)-\partial_{j}m_{d0}\left(x^{C}\right)\partial_{k}\pi_{0}\left(x^{C}\right)\\ \partial_{k}m_{d0}\left(\tilde{x}^{C}\right)\partial_{j}\pi_{0}\left(\tilde{x}^{C}\right)-\partial_{j}m_{d0}\left(\tilde{x}^{C}\right)\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\end{array}\right)=\Psi\left(\begin{array}[]{c}\beta_{d,k}^{C}\\ \beta_{d,j}^{C}\end{array}\right),$ (B.3)
@@
$\Psi=\left(\begin{array}[]{cc}\partial_{j}\pi_{0}\left(x^{C}\right),&-\partial_{k}\pi_{0}\left(x^{C}\right)\\ \partial_{j}\pi_{0}\left(\tilde{x}^{C}\right),&-\partial_{k}\pi_{0}\left(\tilde{x}^{C}\right)\end{array}\right).$
Assumption NL2.(v) ensures that the determinant of $\Psi$ is nonzero, which shows that $\Psi$ is nonsingular. Therefore, equation (B.3) can be solved for $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$ by inverting $\Psi$, thereby identifying $\beta_{d,k}^{C}$ and $\beta_{d,j}^{C}$. Given identification of $\beta_{d,k}^{C}$, we can then identify all other coefficient $\beta_{d,l}^{C}$ in $\beta_{d}^{C}$ by solving (B.2) with the subscript $j$ replaced by $l$, which gives
@@
Appendix C Further discussion on the nonlinearity assumption
@@
When $\dim\left(X^{C}\right)=1$, it follows from (B.1) and the assumption of Theorem C.1 that
@@
The remaining part of the proof is the same as the proof of Theorem 1. ∎
The assumption of Theorem C.1 in place of NL1 requires the univariate function $\pi_{0}$ to be not one-to-one, but imposes no other smoothness or continuity restriction on $\pi_{0}$. Therefore, it doesn’t require $X^{C}$ to be continuously valued. In the extreme case of $X^{C}$ containing only one binary covariate, this assumption will be equivalent to the full irrelevance of $X^{C}$ to the treatment probability, which is a condition suggested by Chamberlain (1986) for identification of the sample selection model. The following two theorems show that Assumption NL1 and the assumption of Theorem C.1 can be extended to the case of multiple continuous covariates.
Theorem C.2.
@@
It follows from equation (B.1) and $\pi_{0}\left(x^{C}\right)=\pi_{0}\left(\tilde{x}^{C}\right)$ that
@@
$\partial_{k}m_{d0}\left(\breve{x}^{C}\right)$ $=$ $\beta_{d,k}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)\partial_{k}\pi_{0}\left(\breve{x}^{C}\right),$ (C.1)
$\partial_{j}m_{d0}\left(\breve{x}^{C}\right)$ $=$ $\beta_{d,j}^{C}+g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)\partial_{j}\pi_{0}\left(\breve{x}^{C}\right).$ (C.2)
Since $\partial_{k}\pi_{0}\left(\breve{x}^{C}\right)\neq 0$ by condition (ii), we can find $g_{d}^{\left(1\right)}\left(\pi_{0}\left(\breve{x}^{C}\right)\right)$ from (C.1) to be
@@
and substitute it into (C.2) to identify $\beta_{d,j}^{C}$ as
@@
Therefore, $\beta_{d}^{C}$ is identified. The remaining part of the proof is the same as the proof of Theorem 1. ∎
Theorem C.3.
@@
The identification of $\beta_{d,k}^{C}$ is analogous to Theorem 1. Then, the identification of all the other coefficients $\beta_{d,j}^{C}$ in $\beta_{d}^{C}$ and the remaining components of the model is the same as in Theorem C.2 and Theorem 1, respectively. ∎
@@
For each $k\in\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$, taking the partial derivative of both sides of (B.1) with respect to $x_{k}^{C}$ gives
@@
which immediately identifies all $\beta_{d,k}^{C}$ in $\beta_{d}^{C}$. The remaining part of the proof is the same as that of Theorem 1. ∎
The above discussion further illustrates the over-identification feature of our strategy. In fact, the strength of identification of our method is determined by the number of nonlinearity points in the data satisfying Assumption NL or Theorems C.1-C.4, given different benchmark values of $X^{D}$.
Appendix D Limited valued outcomes
@@
Assumption D.L (Linear index). Assume that $Y_{d}=l\left(X^{\prime}\gamma_{d},V_{d}\right)$, $d=0,1$, holds for some known link function $l$ and unknown coefficients $\gamma_{d}$, where $V_{d}$ is unobserved error term of $Y_{d}$.
@@
Assumption D.N (Normalization). Decompose $X^{\prime}\gamma_{d}=X^{C\prime}\gamma_{d}^{C}+X^{D\prime}\gamma_{d}^{D}$, where $X^{C}$ and $X^{D}$ consist of covariates that are continuously and discretely distributed, respectively. Assume that $\gamma_{d,1}^{C}$, the first element of $\gamma_{d}^{C}$, equal to 1.
Assumption D.N imposes the convenient normalization that the first continuous covariate has a unit coefficient. This scaling of $\gamma_{d}$ is arbitrary and is innocuous because our focus is on the identification of MTE, which is a difference in the conditional expectations of $l\left(X^{\prime}\gamma_{d},V_{d}\right)$, rather than on separate identification of $\gamma_{d}$ and $l$. Since in the generalized linear model the potential outcomes are not necessarily additive in the error term, the mean independence assumption needs to be strengthened to a stricter distributional independence assumption.
Assumption D.CI (Conditional Independence). Assume that $\left.V_{d}\perp\!\!\!\!\perp X\right|V$ for $d=0,1$, namely, $V_{d}$ is independent of $X$ conditional on $V$, where $V$ is the reduced-form treatment error in equation (3).
Recalling that $V\perp\!\!\!\!\perp X$ by definition, Assumption D.CI is equivalent to the full independence $\left(V_{d},V\right)\perp\!\!\!\!\perp X$ for $d=0,1$, which implies that both the marginal distribution of $V_{d}$ and the copula of $V_{d}$ and $V$ are independent of $X$. Under Assumptions D.L and D.CI, we have
@@
The additive nonseparability of the observables and unobservables constitutes the primary difficulty in identifying MTE in the limited outcome case. Meanwhile, Assumptions D.L and D.CI lead to a double index form of the observable outcome regression functions for each treatment status, that is,
$E\left[\left.Y\cdot 1\left\{D=d\right\}\right|X=x\right]=r_{d}\left(x^{\prime}\gamma_{d},\pi\left(x\right)\right)$ (D.1)
@@
$r_{0}\left(t,p\right)$ $=$ $\int_{p}^{1}E\left[\left.l\left(t,V_{0}\right)\right|V=v\right]dv,$ (D.2)
$r_{1}\left(t,p\right)$ $=$ $\int_{0}^{p}E\left[\left.l\left(t,V_{1}\right)\right|V=v\right]dv.$ (D.3)
Provided that $\pi\left(x\right)$ is a nonlinear index, $\gamma_{d}$ and thus $r_{d}$ can be identified based on functional forms, which ensures identification of MTE since
$\Delta^{\text{MTE}}\left(x,v\right)=\partial_{2}r_{1}\left(x^{\prime}\gamma_{1},v\right)+\partial_{2}r_{0}\left(x^{\prime}\gamma_{0},v\right),$ (D.4)
@@
Assumption D.NL (Non-Linearity). Assume that the functions $\pi_{0}$ and $r_{d}$ are differentiable and denote their partial derivatives with respect to the $k$-th argument as $\partial_{k}\pi_{0}$ and $\partial_{k}r_{d}$, where $\pi_{0}\left(x^{C}\right)=\pi\left(x^{C},0\right)$ and $r_{d}$, $d=0,1$, are defined in (D.1) and (D.1). Assume that there exist two vectors $x^{C}$ and $\tilde{x}^{C}$ on the support of $X^{C}$ and two elements $k$ and $j$ of the set $\left\{1,2,\cdots,\dim\left(X^{C}\right)\right\}$ such that (i) $\partial_{1}r_{d}\left(x^{C\prime}\gamma_{d}^{C},\pi_{0}\left(x^{C}\right)\right)\neq 0$, (ii) $\partial_{1}r_{d}\left(\tilde{x}^{C\prime}\gamma_{d}^{C},\pi_{0}\left(\tilde{x}^{C}\right)\right)\neq 0$, (iii) $\partial_{k}\pi_{0}\left(x^{C}\right)\neq\partial_{1}\pi_{0}\left(x^{C}\right)\gamma_{d,k}^{C}$, (iv) $\partial_{j}\pi_{0}\left(x^{C}\right)\neq\partial_{1}\pi_{0}\left(x^{C}\right)\gamma_{d,j}^{C}$, (v) $\partial_{l}\pi_{0}\left(\tilde{x}^{C}\right)\neq\partial_{1}\pi_{0}\left(\tilde{x}^{C}\right)\gamma_{d,l}^{C}$ for $l=2,\cdots,\dim\left(X^{C}\right)$, and
@@
Assumption D.NL.(i)-(ii) require $r_{d}$ to depend on the linear index, and (iii)-(vi) essentially require some nonlinear variation in $\pi_{0}$ under the scale normalization. As in Assumption NL, it is difficult to construct examples other than $\pi_{0}\left(x^{C}\right)=f\left(x^{C\prime}\gamma\right)$ that violates Assumption D.NL. Finally, a support assumption and an invertibility assumption are imposed for technical reasons.
Assumption D.S (Support). For each $k\in\left\{1,2,\cdots,\dim\left(X^{D}\right)\right\}$, assume for some $x_{k}^{D}\neq 0$ in the support of $X_{k}^{D}$ that there exists $x^{C}\left(k\right)$ in the support of $X^{C}$ such that $\pi\left(x^{C}\left(k\right),x^{Dk}\right)$ is in the support of $\pi_{0}\left(X^{C}\right)$ and that $x^{C}\left(k\right)^{\prime}\gamma_{d}^{C}+x_{k}^{D}\gamma_{d,k}^{D}$ is in the support of $X^{C\prime}\gamma_{d}^{C}$ for $d=0,1$.
Assumption D.I (Invertibility). Assume that the function $r_{d}$ is invertible on its first argument for $d=0,1$.
For the case of a binary outcome, we have $r_{0}\left(t,p\right)=\int_{p}^{1}F_{\left.V_{0}\right|V}\left(\left.t\right|v\right)dv$ and $r_{1}\left(t,p\right)=\int_{0}^{p}F_{\left.V_{1}\right|V}\left(\left.t\right|v\right)dv$, hence a sufficient condition for Assumption D.I to hold is that $V_{d}$ is continuously distributed with support $\mathbb{R}$, conditional on $V$, for $d=0,1$. Interestingly, the same continuous distribution condition also suffices for Assumption D.I in the censored case with $l\left(t,v\right)=\max\left(t-v,0\right)$, where $\partial_{1}r_{0}\left(t,p\right)=\int_{p}^{1}F_{\left.V_{0}\right|V}\left(\left.t\right|v\right)dv$ and $\partial_{1}r_{1}\left(t,p\right)=\int_{0}^{p}F_{\left.V_{1}\right|V}\left(\left.t\right|v\right)dv$. Under the imposed assumptions, the following identification theorem for the generalized linear model follows immediately from Escanciano et al. (2016, Theorems 3.1-3.2).
Theorem D.1.
If Assumptions D.L, D.NL, D.CI, D.S, D.N, and D.I hold, then $\gamma_{d}$ and $r_{d}\left(t,p\right)$ at all points in the support of $X^{\prime}\gamma_{d}$ and $P$ are identified for $d=0,1$.
Having established identification of $\gamma_{d}$ and $r_{d}$, the MTE for limited valued outcomes can be identified by (D.4) and estimated by the semiparametric least squares method of Escanciano et al. (2016). Alternatively, the LIV estimation procedure may be adapted according to